### v1

# TxRay: Agentic Postmortem of Live Blockchain Attacks

Abstract
Decentralized Finance (DeFi) has turned blockchains into financial infrastructure, allowing anyone to trade, lend, and build protocols without intermediaries, but this openness exposes pools of value controlled by code. Within five years, the DeFi ecosystem has lost over B USD to reported exploits. Many exploits arise from permissionless opportunities that any participant can trigger using only public state and standard interfaces, which we call Anyone-Can-Take (ACT) opportunities. Despite on-chain transparency, postmortem analysis remains slow and manual: investigations start from limited evidence, sometimes only a single transaction hash, and must reconstruct the exploit lifecycle by recovering related transactions, contract code, and state dependencies.
We present TxRay, a Large Language Model (LLM) agentic postmortem system that uses tool calls to reconstruct live ACT attacks from limited evidence. Starting from one or more seed transactions, TxRay recovers the exploit lifecycle, derives an evidence-backed root cause, and generates a runnable, self-contained Proof of Concept (PoC) that deterministically reproduces the incident. TxRay self-checks postmortems by encoding incident-specific semantic oracles as executable assertions that the synthesized PoC must satisfy.
To evaluate PoC correctness and quality, we develop PoCEvaluator, an independent agentic execution-and-review evaluator. On 114 incidents from DeFiHackLabs, TxRay produces an expert-aligned root-cause report and an executable PoC for 105 incidents, achieving 92.11 % end-to-end reproduction. Under PoCEvaluator, 98.1 % of TxRay PoCs avoid hard-coding attacker addresses, a +24.8pp lift over DeFiHackLabs, decoupling reproduction from attacker-controlled on-chain state and yielding self-contained, root-cause-consistent reconstructions rather than attacker-contract replays. In a live deployment, TxRay delivers validated root causes in 40 minutes and PoCs in 59 minutes at median latency. TxRay’s oracle-validated PoCs enable attack imitation, improving coverage by 15.6 % and 65.5 % over STING and APE.
1 Introduction
DeFi exposes permissionless on-chain exploit opportunities: finite sequences of transactions that any unprivileged participant can execute using public state and standard interfaces. Some are classical Maximal Extractable Value (MEV) (arbitrage, liquidations, generalized front-/back-runs); others are protocol vulnerabilities that cause damage rather than direct profit (e.g., freezing or permanently locking assets). All are publicly verifiable: once an incident is on-chain, any observer can reconstruct and replay public actions that reproduce the effect without private keys, privileged orderflow, or off-chain agreements. In practice, however, turning a suspicious transaction into a reproducible exploit case relies on brittle heuristics and manual reverse engineering, slowing incident response and hindering systematic measurement. We refer to this class of permissionless, publicly verifiable exploits as ACT opportunities (see Section 3 for the precise definition).
Prior research shows that DeFi exploits are understood retrospectively, with root-cause postmortems arriving hours after the initial loss of funds [zhou2023sok]. The Balancer incident on 3 November 2025 illustrates this lag.111certora blog and balancer postmortem (retrieved November 2025). Malicious transactions began appearing across Ethereum, Arbitrum, Base, Optimism, and Polygon at 07:46 UTC, and automated monitors raised alerts within minutes. A mechanistic root-cause explanation, however, arrived at 10:30 UTC, hours and minutes later, after specialists manually reconstructed cross-chain traces and re-derived the pool invariant math. During that window, structurally similar deployments remained exposed and defenders had to make high-stakes decisions under incomplete understanding. Furthermore, the earliest analyses of the Balancer attack pointed to an incorrect root cause, access control issues, whereas the actual vulnerability stemmed from rounding errors. While these initial explanations circulated, they actively misled the community’s understanding and hindered effective defensive measures for other deployments that shared the same underlying vulnerability.
A second structural problem is that the DeFi ecosystem lacks a dataset that can support scientific evaluation. Public trackers exist, but they do not capture the full exploit mechanism in a reproducible form. DeFiHackLabs runs the most detailed community effort, but its pipeline is manual and resource constrained, so the quality and completeness of PoCs vary by incident. This leads to uneven coverage: some incidents have full traces and carefully engineered PoCs, while others have brief notes. Other trackers from Slowmist, Defillama, De.Fi, Rekt, and BlockSec record incidents independently, but their entries are short summaries and do not include the transaction level steps, invariants, or replayable artifacts needed for verification. The result is a fragmented record that cannot be used to measure coverage, compare detectors, fine-tune AI agents, or study root causes in a consistent way.
We present TxRay, an agentic, LLM-based postmortem system that turns individual incidents into executable cases, addressing both (i) postmortem lag and (ii) the lack of standardized, reusable exploit datasets. Starting from Ethereum Virtual Machine (EVM) seed transaction(s), TxRay first decides whether they reveal an ACT opportunity. If they do, TxRay analyzes the root cause and generates parameter-free PoC tests in Foundry that replay the exploit on a forked blockchain state using a fresh, unprivileged wallet address from scratch, driven by an LLM loop that plans, executes, verifies, and reflects. When the exploit depends on prerequisite actions, TxRay also recovers related transactions (e.g., approvals or liquidity primes), builds the minimal opportunity subgraph, and uses it to guide the postmortem analysis. Conditioned on its inferred root cause, TxRay then generates customized validation oracles, such as profit checks, invariant violations, and consistency checks on the explanation, and requires that all of them pass before accepting a case, ensuring that the final PoC is deterministic, high quality, and faithful to the root cause. TxRay is free to explore within our controlled scope: we provide high-level guidance (e.g., “playbooks”), and TxRay chooses its own path within them. Our primary contributions are:
Agentic postmortem system: We design TxRay, an agentic LLM-based postmortem system that turns individual ACT incidents into structured artifacts: a root-cause report, a parameter-free executable PoC, and a standardized attack lifecycle. From seed transaction(s), TxRay outputs a self-contained bundle of code and metadata that replays on a forked chain and integrates with datasets and tooling, providing the first end-to-end generator of blockchain postmortems with human-analyst quality.
Postmortem correctness and quality evaluation: To measure DeFi PoC correctness and quality, we develop PoCEvaluator, an agentic execution-and-review evaluator: across 120 DeFiHackLabs incidents in a 426-day window (1 Oct 2024–30 Nov 2025, inclusive; after the gpt-5.1 training cutoff), we identify 114 ACT incidents and measure reproduction success (92.11 %) and two error rates: misalignment among ACT-identified cases (3.67 %) and miss rate (4.55 %). On the aligned subset (n=105), TxRay achieves 100% correctness, successfully compiling and passing runtime tests on a pinned on-chain fork. Quality metrics indicate self-contained reconstructions: no attacker-side artifacts appear in 104 / 105 cases; no real attacker addresses appear in 103 / 105 cases; and no exploit-specific constants appear in 53 / 105 cases. Under gpt-5.1 pricing, the estimated median API cost is $3.16 per run for root cause analysis and $1.25 per run for PoC generation.
Live postmortem pipeline: To test TxRay beyond the historical corpus, we deploy our tool in a live setting where a Twitter/X scanner feeds candidate ACT transactions into TxRay during 1 Dec 2025–31 Jan 2026 (62 days, inclusive), continuously processing incidents as they occur. Over this 62-day prospective run, TxRay ingests 24 unique live incident alerts, correctly filters 4 non-ACT cases, and produces 17 aligned root-cause reports on 20 ACT incidents. TxRay achieves a median time-to-PoC of 59 minutes (p95: 105). Among 13 incidents with public expert responses, TxRay beats them (faster with aligned root causes) in 8 cases (62 %), with a median lead time of 192 minutes, directly addressing the postmortem-lag pain point.
LLM-based generalized attack imitation: TxRay is the first LLM-based imitation tool and synthesizes profitable ACT MEV bundles that imitate historical attacks. On the overlap period studied by prior work [qin_usenix23, zhang2023your] (1 Aug 2021–31 Jul 2022; 365 days, inclusive), TxRay finds 5 additional opportunities and improves opportunity coverage by 65.5 % and 15.6 % over two baselines.
Standardized dataset: We generate postmortem artifacts for historical incidents and curate into a single EVM security dataset with normalized taxonomy and labels. We focus on incidents in a 488-day window (1 Oct 2024–31 Jan 2026, inclusive; after the gpt-5.1 training cutoff).222The dataset is continuously updated in production; to finalize this paper, we freeze the snapshot used for all reported results at 31 Jan 2026. This is the most complete dataset of executable DeFi/MEV incidents, cataloging 196 incidents across 9 chains. Every entry corresponds to an accepted postmortem that meets the quality criteria validated in our benchmarking, and the dataset aggregates incidents from DeFiHackLabs, DeFiLlama, De.Fi, Rekt, BlockSec and our live postmortem pipeline.
2 Background
We briefly discuss background on DeFi postmortems.
Blockchains. Public permissionless blockchains in our setting implement a deterministic, account-based state machine. Externally Owned Accounts and contracts share a global state; transactions invoke contract code that can call other contracts, emit logs, and modify balances and storage. On EVM-compatible chains, this behavior is specified by the EVM execution semantics and exposed to off-chain clients via public Remote Procedure Call (RPC) endpoints and archive or tracing infrastructure, including transaction receipts, logs, and detailed execution traces.
Smart contracts, DeFi, and MEV. Smart contracts are programs deployed as on-chain accounts whose code and state are executed by the EVM. DeFi protocols are collections of such contracts that implement financial logic, swaps, lending, liquidations, and structured products. Their composability enables complex cross-protocol interactions but also creates value-redistribution opportunities commonly referred to as MEV: sequences of transactions that extract value by reordering, inserting, or copying interactions (e.g., arbitrage, liquidations, sandwiches, and generalized back-/front-runs). Prior work surveys DeFi attacks and MEV behaviors in detail [zhou2023sok, qin_usenix23].
DeFi attacks, IDS, and postmortems. DeFi attacks are sequences of on-chain actions that violate protocol safety or liveness properties, e.g., by exploiting logic bugs, misconfigured parameters, or adversarial interactions between protocols. Many attacks are permissionless and publicly reproducible; in Section 3 we formalize ACT as the subset that an unprivileged adversary can realize using only public information and standard interfaces. When an exploit or anomalous behavior occurs, practice typically unfolds in two stages. First, intrusion-detection systems (IDS) (e.g., protocol-specific detectors, on-chain monitoring platforms, and incident dashboards) signal that an incident is in progress. Research and industry efforts focus heavily on this detection layer, from LLM-based anomaly detection [gai2023blockchain] to commercial systems such as BlockSec Phalcon [blocksec_dashboard], Hypernative [hypernative], and similar real-time monitoring services. Second, analysts perform postmortem reconstruction: recovering the precise sequence of on-chain actions, identifying the underlying vulnerability, and producing an executable PoC. To the best of our knowledge, this stage remains manual, leaving teams and the ecosystem to act under incomplete understanding: incidents may be misclassified as benign or non-exploitable, decisions may be wrong, and losses may continue until later analyses recover the true exploit mechanism and affected surface (e.g., the Balancer precision-loss exploit [balancer_attack_weilinli]).
Generalized attack imitation. The visibility of pending transactions enables imitation and MEV-time counterattack systems, where an adversary (or defender) observes a transaction and synthesizes a semantically similar bundle to redirect value or preempt theft. Prior systems such as APE [qin_usenix23] and STING [zhang2023your] automate this process. STING achieves greater coverage than APE by incorporating heuristics (e.g., fund provenance, contract verification status) to identify adversarial contracts, whereas APE relies on program execution traces and taint analysis. We position TxRay relative to them in Section 7. While TxRay targets postmortem reconstruction, its PoCs can also be reused for generalized imitation.
3 Models
In the following, we formalize the system and threat models.
3.1 System model
We assume an EVM-compatible blockchain with (i) a canonical chain chosen by the underlying consensus protocol; (ii) a bounded reorganization depth that is short relative to incident analysis, so that confirmed blocks are stable; and (iii) archive or tracing infrastructure that exposes the same on-chain data to all participants, including canonical state, receipts, logs, and execution traces (e.g., call graphs and per-opcode effects). When contract source code and Application Binary Interfaces are published via public explorers (e.g., Etherscan-style services), we assume they are accurate and publicly visible. For unverified contracts, we assume only bytecode is available and ground analysis in call interfaces, traces, and state diffs; public disassembly or decompilation can aid this process but is not assumed correct or complete. In both cases, we assume public RPC endpoints implement canonical state queries and local transaction simulation, enabling any party to reconstruct historical state at a chosen block and deterministically replay transactions from that fork. These assumptions restrict our deployment to chains or infrastructures where trace, replay, explorer, and RPC APIs are available. We do not require any non-public data or validator internals, and we abstract away off-chain channels (e.g., phishing or misleading wallet UIs), considering only on-chain execution semantics and publicly observable infrastructure.
3.2 Threat model
ACT opportunities. We use “anyone-can-take” (ACT) for permissionless on-chain exploits an unprivileged adversary can realize using public data and standard interfaces. We exclude phishing and social-engineering attacks (e.g., address poisoning [guan2024characterizing, ye2024interface, tsuchiya2025blockchain]) whose success depends on victim mistakes and is probabilistic. Accordingly, when we refer to a “victim transaction” below, we mean an intended on-chain interaction (e.g., a trade, borrow, or liquidation) rather than a transaction induced by off-chain deception (phishing). Formally, we define an ACT opportunity as follows. Fix a chain and block height , and let denote the pre-state reconstructible from canonical on-chain data (via RPC, logs, and traces) and public contract metadata (e.g., verified source code and ABIs on explorers, when available). Without published metadata, is derived from bytecode and traces. An ACT opportunity exists at if there exist and a transaction sequence that satisfy all of the following:
1.
for each , is a transaction that the unprivileged adversary can include under ’s standard submission and inclusion rules, and it is either:
an attacker-crafted transaction, constructed and signed by the adversary using only data visible at or before under the system model (i.e., public on-chain data, traces, and explorer metadata); or
a victim transaction constructed and signed by a third party that is observable to the adversary via public infrastructure (e.g., the mempool or builder/relay feeds), and whose contents the adversary does not modify (excluding phishing-induced transfers);
2.
executing from under the EVM transition function deterministically yields a post-state in which at least one of the following exploit predicates holds:
a profit predicate, where the adversary’s net portfolio value in a fixed reference asset increases after fees, i.e.,
where is the adversary address, maps a state and address to total value in a reference asset (e.g., ETH or USD), and is the total gas paid by ; or
a non-monetary predicate, where there exists a deterministic, publicly checkable safety/liveness predicate such that (e.g., an asset becomes permanently unusable, a collateralization invariant is broken, or a governance threshold is crossed).
Adversary model. The adversary is a permissionless on-chain actor that can realize ACT opportunities. The adversary controls one or more unprivileged EOAs, can deploy contracts, and observes the information available under our system model (canonical state, logs, traces, explorer metadata, and transactions visible via the public mempool or builder/relay feeds). The adversary can submit transactions or ordered bundles for inclusion, choose gas prices and inclusion fees, and run off-chain computation (e.g., local simulation and search). We do not grant the adversary cryptographic breaks, victim private keys, privileged control over consensus, or visibility into non-included private orderflow; private bundles become observable once included on-chain. We also do not model adversarial influence over off-chain channels (e.g., phishing or misleading wallet UIs) and thus exclude probabilistic attacks that rely on inducing victim mistakes.
Scope disclaimer. Our threat model includes all opportunities on EVM-compatible chains that satisfy our ACT definition, whether via the profit predicate or the non-monetary predicate, and whether labeled as MEV or “attacks”.
4 TxRay Design
TxRay turns seed transaction(s) into a validated postmortem. We assume an external alerting pipeline provides the chain id(s) and transaction hash(es). TxRay outputs (i) an evidence-backed root cause analysis and (ii) a self-contained Foundry PoC that reproduces the incident on a forked chain state, together with structured validation artifacts used for quality benchmarking. Figure 2 summarizes the workflow: a shared orchestrator coordinates six specialist subagents333Implementation detail: the orchestrator interacts with each subagent through a tool-call style interface, which simplifies logging and resource control; each subagent can call external tools/APIs (cf. Figure 2). over a shared session workspace, iterating with validator feedback under an explicit turn budget.
4.1 Shared Components
All stages share (i) a central orchestrator and (ii) a workspace.
Orchestrator controls all stages (Figure 2): it instantiates a workflow plan, executes specialist subagents, and routes control based on their structured outputs. After each step, it checks for expected on-disk artifacts (e.g., traces, reports, and synthesized tests) and validates them (e.g., JSON schema conformance). When a validator/challenger rejects, it returns structured feedback (e.g., missing evidence or lifecycle gaps), which the orchestrator maps to re-analysis or re-collection; otherwise the orchestrator advances or aborts when the seed is not an ACT opportunity.
Workspace shares artifacts across stages, so outputs produced by earlier steps are visible to later ones. Specialist tools read inputs (e.g., collected transactions, traces, and contract artifacts) and write outputs as artifacts. The orchestrator persists these artifacts and uses them as the interface between tools, supporting reruns and targeted re-collection when validation indicates missing lifecycle coverage.
4.2 Root Cause Analysis
Root cause analysis (components (1)-(3) in Figure 2) maps seed transaction(s) to (i) a lifecycle transaction set and (ii) a root-cause report grounded in on-chain evidence. It iterates evidence collection, analysis, and validation until the validator accepts or the budget is exhausted, and aborts when the seed transaction(s) do not correspond to an ACT opportunity.
(1) On-chain data collector given the seed transaction(s), gathers evidence available to a public observer and materializes three seed artifacts: (i) transaction metadata (e.g., sender/receiver, calldata, block), (ii) an execution trace, and (iii) balance and storage diffs derived from receipts/logs and state-diff tracing. In our implementation, the collector queries explorer APIs (e.g., Etherscan v2), public RPC endpoints (e.g., eth_getTransactionReceipt and debug_traceTransaction with prestateTracer), and Foundry/cast; when contracts are unverified, it falls back to bytecode-level artifacts (e.g., disassembly and optional decompilation with tools such as Heimdall [heimdall_rs]). The orchestrator then hands these artifacts to the root cause analyzer. When the analyzer or challenger identifies missing evidence (e.g., setup or funding transactions referenced by the draft), the orchestrator forwards targeted data requests (e.g., additional traces or contract artifacts) and the collector appends the results to the shared workspace.
(2) Root cause analyzer synthesizes artifacts into a root-cause report draft that identifies victim contracts and attacker-controlled roles, localizes the code path or invariant violation, and explains the exploit mechanism. Because ACT opportunities can span multiple transactions, it expands the seed transaction(s) into a set of related transactions covering the lifecycle (e.g., initial funding, approvals or liquidity primes, attack-contract deployment, exploit execution, and profit extraction) and iteratively requests missing on-chain evidence from the data collector. To remain evidence-first, each claim must be grounded in traces, state diffs, or verified source/bytecode. After deciding whether the seed is an ACT opportunity, the analyzer writes the (ACT or non-ACT) root-cause report draft and hands it to the root cause challenger for verification.
(3) Root cause challenger checks the report draft against collected artifacts. It rejects drafts that omit lifecycle steps or roles, contradict traces/state diffs, or contain undetermined claims, and it requires an end-to-end causal account that states the violated invariant and the code-level breakpoint. On rejection, it returns structured feedback and missing-evidence requests; the orchestrator routes to re-analysis, re-collection, or seed-to-lifecycle expansion (Appendix A).
4.3 PoC Generation
After validating the root cause, TxRay synthesizes an executable PoC through three specialist subagents (components (4)–(6) in Figure 2) coordinated by the orchestrator.
(4) Poc oracle generator translates the validated root cause into a semantic oracle definition. The definition has three parts: (i) variables naming entities (victim contracts and assets by address, attacker roles as logical identifiers with null addresses); (ii) pre_check encoding sanity checks (e.g., reward pool funded, non-zero stake); and (iii) oracle_constraints enumerating hard and soft constraints (category, natural-language description, and pseudocode) for Foundry tests. Hard constraints encode non-negotiable properties: permission/ownership changes, critical logic or state invariants, revert/guard semantics, and asset identity (e.g., profit must be realized in the same token type as the incident). Soft constraints encode economic directionality with tolerance (e.g., attacker profit and victim depletion) using inequalities with explicit thresholds and rationale (e.g., profit within tolerance of the incident’s observed profit, rather than only “strictly positive profit”). The oracle generator omits execution details that are not part of the mechanism (timestamps, gas accounting, incidental counter updates) and does not depend on attacker-side identities or artifacts from the incident. This oracle definition supports semantic checking without bit-for-bit replay of the original transactions.
(5) Poc reproducer implements a self-contained Foundry project that forks chain state at the incident block, re-creates funding, setup, and contract deployment, executes the lifecycle using fresh attacker-role addresses, and encodes the oracle definition as assertions in the Foundry tests. It avoids using the real attacker EOAs, attacker-deployed contract addresses, incident calldata, or attacker-side artifacts.
(6) Poc validator runs Foundry tests on a fork and checks (i) oracle satisfaction (without mocked contract behavior) and (ii) a rubric (readability, absence of magic numbers, self-containment, and lifecycle documentation). On failure, it returns rejection reasons and fixes; the orchestrator reruns the reproducer. On success, TxRay outputs the report, the Foundry PoC, and structured validation results that we aggregate into a postmortem quality score.
4.4 Implementation Details
Implementation. TxRay is implemented in 2,255 lines of Python on top of the OpenAI Agents SDK (v0.6.1). The orchestrator calls the OpenAI API; the six specialist subagents (components (1)–(6) in Figure 2) run via the Codex CLI (v0.63.0). Each specialist subagent runs as a stateful Codex conversation (invoked via codex exec and resumed by conversation id) with a system-prompt template instantiated with session paths. We run gpt-5.1 with medium thinking budget (knowledge cutoff: 30 Sep 2024) for both the orchestrator and Codex-backed tools, and cap each stage at max_turns=60. We disable Codex web search and browsing in config.toml.
Runtime environment. Our experiments run on an Ubuntu 22.04.5 server with an Intel i9-12900KS (16 cores / 24 hardware threads), 62 GiB RAM, and a 699 GiB NVMe system volume. Each run uses a session directory containing a .env file with credentials: (i) OPENAI KEY; (ii) ETHERSCAN KEY; and (iii) QUICKNODE KEY for archive RPC endpoints. At session start, the orchestrator copies JSON schemas and helper scripts, resolves chain-specific RPC URLs from a chainid-to-template map, and persists artifacts (traces, source or bytecode, JSON reports, Foundry projects, and logs). TxRay supports 22 EVM chains (Appendix C), requiring (i) archive+debug_traceTransaction RPC access (QuickNode), (ii) Etherscan v2 explorer APIs (account txlists, verified sources when available), and (iii) Foundry forked execution for PoC validation; when sources are unavailable, we use Heimdall as a bytecode-level fallback.
4.5 Walkthrough
Appendix A provides a compact end-to-end walkthrough with minimal iteration, giving intuition for the artifact flow and the orchestrator’s turn-by-turn control loop. In this case, our incident monitor extracts the seed transaction hash from a Twitter/X alert at 09:06 UTC on 1 Jan 2026; TxRay produces a validated root-cause report in 38 min and a validated PoC in 47 min from the alert. We highlight the Valinity walkthrough (Appendix B), which stress-tests TxRay under unverified attacker-side components and protocol-specific economic preconditions. The seed transaction calls an unverified contract, and the exploit hinges on wiring and preconditions not explicit in the seed trace. TxRay reconstructs an evidence-backed ACT narrative in which a flash-loan-driven swap sequence manipulates prices, triggers acquireByLTVDisparity() to mint VY and inflate the PAXG cap, and then reuses VY as collateral to open loans that unwrap and extract ETH-equivalent from WETH9 reserves. The following key takeaways summarize the main challenges in this walkthrough:
(i) Filling information gaps under partial observability. Starting from a seed transaction, the full exploit context must be recovered, such as upgradeable wiring (proxyimplementation), registrar lookups, and per-asset configurations, and verified at the incident block. TxRay performs this autonomously by reconstructing , the on-chain pre-state in our ACT definition (Section 3.2), using targeted storage reads and state diffs rather than relying on inferred contract graphs. It then produces an ACT-compatible explanation of the transition . Result: In the Valinity case, closing these gaps required only four storage-slot queries and a pre-state diff trace for the seed transaction, successfully recovering the registrar wiring and per-asset configuration needed for an ACT-complete explanation in our Valinity walkthrough (Appendix B, Turns 3–6).
(ii) Reasoning with imperfect and low-level artifacts. When attacker-side code is unavailable, analysts often rely on decompilation/disassembly (e.g., proprietary tools such as DeDaub) and debuggers (e.g., BlockSec Phalcon) that lift raw traces and state diffs into higher-level views. TxRay instead uses open-source bytecode tooling (less accurate than proprietary decompilers) and low-level RPC/explorer APIs, so it must lift evidence into higher-level claims while staying robust to imperfect decompilation. It grounds feasibility in txlists and opcode-level evidence (traces, disassembly, diffs), treating decompiler output as a hint and revising hypotheses when trace evidence contradicts it. Result: TxRay proceeded without verified attacker code by combining address-history queries, opcode-level traces (including a 14-transaction window around the incident block), and raw router bytecode/disassembly to justify unprivileged feasibility (Appendix B, Turns 2 and 6).
(iii) Quality-driven iteration to reach reproducible artifacts. Human postmortems often stop at a plausible narrative and a “good-enough” reproducer (e.g., a shallow replay with hard-coded parameters). Our benchmark standard is higher: the final case must be executable, self-contained, parameter-free, and satisfy semantic oracles encoding the claimed mechanism. We enforce this with validator-driven rejection of non-conforming drafts; in our run, the reproducer required three attempts (two rounds of rejection and revision) before producing a self-contained Foundry PoC that re-creates the on-chain conditions (e.g., via a USDCPAXG Uniswap V3 swap) and satisfies the semantic oracles (Appendix B, Turns 7–9).
Overall, the Valinity run required six analyzer iterations and five data-collection iterations, during which the data collector executed 18 evidence fetches. The PoC validator rejected two drafts before accepting the final self-contained exploit. End to end, TxRay produced a validated root-cause report in ~64 min and a validated PoC in ~92 min.
5 Postmortem Quality Benchmarking
We benchmark TxRay against the community baseline, DeFiHackLabs, and ask four practitioner-facing questions:
RQ1 (Fidelity). Given seed transaction(s) from DeFiHackLabs, in what fraction of incidents does TxRay (i) identify an ACT opportunity and (ii) produce a root-cause report aligned with expert ground truth?
RQ2 (Reproducibility). For aligned cases, in what fraction does TxRay synthesize an executable, self-contained, oracle-satisfying PoC, and how does it compare to DeFiHackLabs under PoCEvaluator?
RQ3 (Reliability). How effective are TxRay’s verification stages (root cause challenger and poc validator) at distinguishing invalid root causes and invalid PoCs, and what failure modes and rejection reasons dominate?
RQ4 (Practicality). What are latency (wall-clock runtime), token usage, and estimated cost per incident?
Dataset. We benchmark on 120 public incidents from DeFiHackLabs dated between 1 Oct 2024 and 30 Nov 2025 (426 days, inclusive). We exclude six non-ACT incidents, five rug pulls (BUBAI, VRug, Roar, IRYSAI, and YziAIToken) and one supply-chain attack (Bybit), and extract the seed transaction hashes for the remaining 114 incidents.
Summary of results. On 114 ACT incidents, TxRay identifies 109 as ACT opportunities and produces expert-aligned root-cause reports for 105; this corresponds to a 3.67 % misalignment rate among ACT-identified cases and a 4.55 % miss rate. TxRay synthesizes executable, self-contained PoCs for 105 incidents (92.11 %), with median times of 29.39 min (root cause analysis) and 10.65 min (PoC generation)..
5.1 Root Cause Analysis Effectiveness
We run TxRay on the 114 selected ACT incidents, executing one run per incident. After completion, two professional blockchain security researchers independently review TxRay’s final reports and the generated codebases, and then cross-validate their assessments. We use two binary labels: (i) ACT-identified indicates whether TxRay identifies an ACT opportunity from the seed transaction(s), and (ii) root-cause match indicates whether the inferred exploit mechanism matches the ground truth. Table 4 reports these labels (columns ACT and RC) and the derived outcomes below.
Aligned: the incident is ACT-identified and has a root-cause match, i.e., TxRay identifies an ACT opportunity and matches the ground-truth exploit mechanism.
Misaligned: the incident is ACT-identified but does not have a root-cause match, i.e., TxRay identifies an ACT opportunity but reports a misaligned root cause.
Missed: the incident is not ACT-identified (root-cause match undefined), i.e., given the seed transaction(s) TxRay either concludes non-ACT or cannot construct an evidence-backed ACT opportunity.
| Description | |
| C1 | Does the PoC compile under Foundry (e.g., forge test builds)? |
| C2 | Does the PoC run without reverts and satisfy the intended oracles? |
| C3 | Does the PoC run on a pinned on-chain fork (not only local mocks)? |
| Q1 | Does the PoC avoid attacker-side artifacts (e.g., attacker-deployed helper contracts), re-implementing the attack from scratch? |
| Q2 | Does the PoC avoid real attacker addresses (use deterministic fresh test addresses instead, e.g., makeAddr)? |
| Q3 | Does the PoC avoid attacker-specific hard-coded values (e.g., calldata, parameters, amounts, crafted constants)? |
| Q4 | Does the PoC assert success predicates (e.g., asset deltas, state changes, invariant breaks), not only completion? |
| Q5 | Does the PoC include comments for non-obvious calls and parameters? |
| Q6 | Does the PoC label key addresses and roles (e.g., Attacker, Victim)? |
Results. On the 114 selected ACT incidents, TxRay identifies 109 cases as ACT opportunities and produces corresponding root-cause reports. Among the ACT-identified cases, 105 root-cause reports match the ground truth. Under the mapping above, this yields a 3.67 % misalignment rate among ACT-identified cases (4/(105+4)) and a 4.55 % miss rate (5/(105+5)). Table 4 reports the per-incident outcomes.
Failure Case Analysis. We investigate the 9 cases where TxRay does not produce an expert-aligned root cause and attribute them to four limitations:
(i) Semantic decoding of low-level inputs (2/9, 22.2 %): when the on-chain data collector cannot obtain a complete and accurate ABI (Figure 2, component (1)), the root cause analyzer (component (2)) must decode binary-encoded calldata from raw bytes, which limits recovery of semantic fields (OneInchFusionV1SettlementHack, LeverageSIR).
(ii) Seed-to-lifecycle reconstruction (2/9, 22.2 %): when the dataset seed transaction(s) capture profit extraction or downstream effects rather than the vulnerability trigger, the on-chain data collector (Figure 2, component (1)) fails to retrieve prerequisite and trigger transactions needed to reconstruct the full lifecycle, leaving root-cause localization under partial evidence (BalancerV2, DeltaPrime).
(iii) Protocol-intent root-cause reasoning (2/9, 22.2 %): even with sufficient on-chain context collected (Figure 2, components (1)–(3)), TxRay fails to connect code-level behavior to the exploit mechanism, yielding a misaligned explanation; this reflects limitations of the underlying language model’s protocol-level reasoning (Laundromat, Corkprotocol).
(iv) ACT boundary and economic-mechanism confusion (3/9, 33.3 %): despite an explicit ACT definition, TxRay misattributes profit to protocol-intended value flows or market-dependent mechanics (e.g., fee-on-swap), treating them as the trigger and diverging from our ACT framing and the ground-truth mechanism (P719Token, YBToken, WXC_Token).
5.2 PoC Correctness and Quality Evaluation
We compare the PoCs synthesized by TxRay against the DeFiHackLabs baseline along two dimensions: (i) Correctness measures functional executability (C1–C3 in Table 1): the PoC compiles under Foundry, executes the intended exploit flow without unexpected errors/reverts, and runs on a deterministic on-chain fork pinned to a specific block height (or block hash). (ii) Quality measures whether the PoC is a self-contained and readable reconstruction of the exploit (Q1–Q6): it avoids attacker-side artifacts, real attacker addresses, and attacker-designed hard-coded constants; it encodes explicit success predicates as assertions; and it improves readability via comments and address labels.
PoCEvaluator Design. We implement PoCEvaluator to evaluate TxRay and DeFiHackLabs under a shared checklist (Table 1). Specifically, we follow CodeVisionary [wang2025codevisionaryagentbasedframeworkevaluating], which evaluates generated code via independent execution-based evaluation followed by multi-round negotiation and reports improved correlation with human judgments (+0.217 Pearson, +0.163 Spearman, +0.141 Kendall Tau over the best baseline on a 363-sample benchmark). PoCEvaluator instantiates this protocol for DeFi PoCs (Figure 3) in two stages. In stage (i) Agentic Independent Evaluation, it takes a PoC project directory and dispatches sub-evaluator agents to run forge test, inspect code and logs, and record execution artifacts (e.g., compilation status and pass/fail signals); sub-evaluators may also query on-chain context via Etherscan APIs, QuickNode RPC, and cast. In stage (ii) Multi-round Negotiation, an aggregator agent aggregates reports, isolates disagreements, and prompts sub-evaluators to maintain or revise decisions based on evidence until convergence, producing a consolidated report with decision histories (Appendix F).
Correctness Evaluation Result. We run PoCEvaluator with independent sub-evaluator agents. End-to-end on the 114 ACT incidents, TxRay produces runnable PoCs for 105 cases, achieving an end-to-end PoC reproduction rate of 92.11 % (105/114). For incidents without a root-cause–aligned PoC (i.e., misaligned or missed under Table 4), downstream execution outcomes are not meaningful for cross-method comparison, as failures stem from upstream misalignment rather than PoC executability; we therefore mark the downstream correctness/quality metrics as invalid (shown as “–”) and exclude them from the head-to-head comparison. On the aligned subset (n=105), all TxRay PoCs compile, execute successfully, and use a pinned on-chain fork (C1–C3: 105/105 each; Figure 4). The DeFiHackLabs baseline matches on compilation and forked execution (C1/C3: 105/105 each) but has two run failures (C2: 103/105): 2025-08/EverValueCoin_exp and 2025-01/Paribus_exp; we manually re-ran both cases and likewise could not execute them successfully, observing EvmError: revert and Transfer failed, respectively.
Quality Evaluation Result. On the aligned subset (n=105), Q1–Q3 measure whether a PoC reflects the inferred root-cause model and reproduces the strategy without relying on attacker-side artifacts, attacker account state, or attacker-designed constants (e.g., hard-coded calldata or exploit-specific numeric parameters). Under these metrics, TxRay passes Q1 in 104/105 cases (vs 102/105 for DeFiHackLabs), Q2 in 103/105 (vs 77/105), and Q3 in 53/105 (vs 14/105; Figure 4). Q4 (has_success_predicate) checks whether the PoC encodes validation oracles as assertions; TxRay passes Q4 in 105/105 aligned cases, and DeFiHackLabs passes Q4 in 3/105. Q5 (has_explanatory_comments) and Q6 (uses_address_labels) measure readability; TxRay passes Q5 in 93/105 cases (vs 63/105) and Q6 in 105/105 (vs 94/105).
Manual Review. To validate that PoCEvaluator aligns with expert judgment, two authors independently inspect a time-stratified sample of 20 incidents from the aligned subset (20/105), manually assessing PoC correctness and quality against the metrics. We find full agreement between the manual assessments and PoCEvaluator on all sampled cases.
5.3 Challenger and Validator Effectiveness
TxRay employs a reject–refine loop for root cause analysis and PoC generation: iteratively rejecting and refining non-conforming intermediate artifacts with structured feedback. Beyond end-to-end success, we evaluate its guardrails by analyzing rejection rates and reasons from the root cause challenger and poc validator. In the root cause analysis stage, the challenger issues a median of 1 rejects (p90: 2; max: 4). Figure 6 shows that unknown content (42 %) and speculative language (43 %) dominate. In practice, these rejects push the root cause analyzer to tighten grounding by fetching and citing additional evidence (e.g., state/balance diffs, proxyimplementation resolution, and missing transaction context), reducing hallucination and improving determinism. More specific reasons enforce targeted checks: missing onchain traces (7 %) forces the analyzer to justify each causal step from executed call traces (rather than inferring solely from diffs or code), and incomplete ACT lifecycle requires an end-to-end reconstruction of the attacker’s preparationtriggersuccess-predicate path. In the PoC generation stage, reject counts range from 0 to 4 (p90: 2). The dominant rejects are uses attacker-designed values (24 %) and uses attacker contract (24 %), which enforce self-containment by disallowing incident-specific hard-coded calldata/parameters and attacker-deployed contract addresses. These constraints force the poc reproducer to re-derive the exploit from public on-chain state and re-synthesize minimal artifacts. We also observe oracle validation failed (17 %): during forge test on a fork, the PoC fails one or more hard/soft validation oracles produced by the poc oracle generator.
5.4 Cost and Performance Analysis
We collect and aggregate TxRay’s runtime and token usage while executing the DeFiHackLabs benchmark.
Token Usage Analysis. Figure 5 (left) breaks token usage into cached input, uncached input, and output tokens. Uncached input tokens are newly processed prompts, cached input tokens are prompt tokens served from caching (reused request prefixes), and output tokens are model-generated. The cached-input category is large because agentic execution repeatedly resends an invariant prompt prefix across many tool-calling turns (e.g., system instructions, tool schemas, accumulated artifacts), which becomes cacheable and is billed at a discounted rate. Across runs, cached input accounts for the largest share: root cause analysis consumes a median of 7.94M cached input tokens (IQR: 6.21M–10.70M) versus 0.50M uncached input tokens (IQR: 0.38M–0.60M), while producing 0.158M output tokens (IQR: 0.118M–0.191M). PoC generation follows the same pattern, with medians of 3.60M cached input, 0.168M uncached input, and 0.0566M output tokens.
Cost Estimation. We estimate per-run cost under a direct API setting, since our Codex subscription does not expose per-request charges. Using gpt-5.1 pricing during the experiment ($1.25/M uncached input; $0.125/M cached input; $10.00/M output), we compute . Figure 5 (middle) reports the resulting cost distributions. Root cause analysis has a median estimated cost of $3.16 per run (IQR: $2.57–$3.93), with a maximum of $6.32. PoC generation has a median estimated cost of $1.25 (IQR: $0.92–$1.93), with a maximum of $9.38; For visualization, the range is capped at $8, excluding a single outlier.
Execution Time Analysis. Figure 5 (right) shows the distribution of end-to-end session time for the root cause analysis stage and the PoC generation stage. Root cause analysis has a median session time of 29.39 minutes (IQR: 21.07–36.78), and PoC generation has a median of 10.65 minutes (IQR: 7.13–20.52). Figure 7 breaks down end-to-end latency by pipeline component. The heavy-tailed components correspond to the model-facing reasoning steps, with median per-incident time of 18.6 minutes for the root cause analyzer and 7.0 minutes for the poc reproducer; the remaining tools (data collection, challenging, oracle generation, and validation) contribute smaller and more stable overhead. Orchestrator time is the residual coordination overhead after subtracting tool execution time. Figure 8 reports the latency distribution of calls to the most frequently invoked downstream tools/commands/scripts in TxRay’s pipeline. Most single calls complete quickly (top; median: 1.2 seconds; p95: 8.8 seconds), but the pipeline issues many such calls per incident; the per-incident time-share plot (bottom) shows how their cumulative overhead can occupy a material share of wall-clock time. While most individual calls are fast, repeated RPC/API queries and command executions can accumulate into a material share of time, highlighting optimization opportunities in the surrounding toolchain infrastructure (e.g., caching/batching and parallelization)
| Incident Name | Incident Provider | ACT Incident | Estimated Loss (USD) | Incident Post Time (UTC) | Monitor Finished | Root Cause Finished | PoC Finished | Expert Response | TxRay | Compact | Monolith | |||||||
| RC | PoC | RC | PoC | RC | PoC | |||||||||||||
| DMi | TenArmorAlert | $124.4K | 25-12-09 02:34 |
AL | MS | – | AL | |||||||
| HTC Token | TikkalaResearch | $45K | 25-12-09 20:14 |
AL | MS | – | AL | |||||||
| WADJET | TenArmorAlert | $97.5K | 25-12-15 02:45 |
AL | MS | – | AL | |||||||
| Dragun69 | TenArmorAlert | $87.4K | 25-12-22 02:10 |
MA | – | MA | – | MA | – | |||||
| 0x530_Transfer | CertiKAlert | $2.3M | 25-12-23 08:45 |
– | – | NA | – | NA | – | NA | – | |||||
| MorningStar | nn0b0dyyy | $12.9K | 25-12-25 06:11 |
– | AL | AL | AL | ||||||||
| DoomCat | TikkalaResearch | $5K | 25-12-30 17:59 |
– | AL | MS | – | MS | – | ||||||
| PRXVTai | CertiKAlert | $97K | 26-01-01 09:06 |
AL | MS | – | AL | |||||||
| Valinity | TenArmorAlert | $18K | 26-01-04 02:46 |
AL | MS | – | AL | |||||||
| Unverified_0x46e1 | TikkalaResearch | $60K | 26-01-05 17:05 |
– | AL | AL | MS | – | |||||||
| IPOR | CertiKAlert | $330K | 26-01-07 02:31 |
MS | – | MS | – | MS | – | |||||
| UXLINK | getblock_en | $23M | 26-01-07 07:32 |
– | – | NA | – | NA | – | NA | – | |||||
| Truebitprotocol | CertiKAlert | $26.4M | 26-01-09 04:18 |
– | MA | – | MA | – | MA | – | |||||
| Unverified_0xf7ca | TenArmorAlert | $394.7K | 26-01-12 02:07 |
– | AL | AL | AL | ||||||||
| Metaverse Token | TenArmorAlert | $37K | 26-01-12 08:28 |
AL | MA | – | AL | |||||||
| YO Protocol | GoPlusZH | $3.7M | 26-01-14 00:59 |
– | – | NA | – | NA | – | NA | – | |||||
| SynapLogic | TenArmorAlert | $88K | 26-01-20 02:06 |
AL | MA | – | MA | – | ||||||
| Makina Finance | TenArmorAlert | $4.2M | 26-01-20 03:54 |
AL | MS | – | MA | – | ||||||
| AiPay | TenArmorAlert | $30.1K | 26-01-23 09:45 |
AL | MA | – | MA | – | ||||||
| SwapNet (base) | CertiKAlert | $13.3M | 26-01-25 21:34 |
AL | AL | AL | ||||||||
| SwapNet (eth&bsc) | TenArmorAlert | $16.8M | 26-01-26 02:52 |
– | AL | AL | AL | ||||||||
| PGNLZ | TenArmorAlert | $100.9K | 26-01-28 03:00 |
– | AL | MA | – | AL | |||||||
| XPlayer | CertiKAlert | $717k | 26-01-28 15:26 |
AL | MS | – | MS | – | ||||||
| Fishing_0xd674 | GoPlusZH | $12.4M | 26-01-31 06:18 |
– | – | NA | – | NA | – | NA | – | |||||
6 Live Postmortem Pipeline
In this section, we evaluate TxRay in a prospective live deployment to assess its effectiveness on previously unseen incidents under real-world latency constraints. We define upstream delay as the time from the first public alert post to our monitor forwarding the seed transaction hash(es) and chain id(s) to TxRay (postmonitor). Since building a transaction-level Intrusion-Detection System (IDS) is not a focus of this paper, we implement a Twitter/X-based Incident Monitor that extracts candidate incident transaction hash(es) mentioned in public posts. We run the end-to-end pipeline continuously during 1 Dec 2025–31 Jan 2026 (62 days). Each incident is run once (no cherry-picking).
Postmortem Pipeline Implementation. Incident Monitor subscribes to a set of Twitter/X accounts (Appendix D). We seed this list with (i) incident announcement accounts and analysts referenced in DeFiHackLabs and (ii) additional accounts found via keyword search and manual screening. When a subscribed account posts, the monitor extracts txhash(es), resolves chain id(s), de-duplicates incidents, and uses gpt-oss-120b to filter for incident-related posts. For posts that pass, it forwards only txhash(es) and chain id(s) to TxRay (no narrative context), so TxRay reasons from on-chain evidence. We use Incident Monitor as an incident source and do not evaluate its coverage or precision. We seek to answer the following research questions:
RQ5 (Prospective Accuracy). Given incident seeds forwarded by our monitor (postmonitor), how often does TxRay (i) identify an ACT opportunity and (ii) produce a root-cause match, while filtering non-ACT alerts?
RQ6 (Time-to-Postmortem). For ACT incidents whose first alert does not include a root-cause explanation, what are TxRay’s time-to-root-cause and time-to-PoC, and how do they compare to the first correct public expert analysis?
RQ7 (Ablation Study). How do TxRay’s prompt and workflow design choices affect prospective performance under a live incident stream?
6.1 Live Postmortem Effectiveness
Table 2 reports our live deployment results. The Incident Monitor forwarded 24 unique candidate DeFi incident posts (each with seed transaction hash(es)) to TxRay for end-to-end postmortem. The Incident Provider column reports the first account that supplied the incident transaction hash(es), and ACT Incident indicates whether the incident is an ACT opportunity under our threat model. Manual Validation reports expert review of TxRay outputs: Aligned, MisAligned, and Missed follow Section 5.1; Non-ACT indicates a non-ACT seed transaction that TxRay rejects. TxRay correctly rejects all 4 non-ACT seed transactions. On the 20 ACT incidents, we observe 17 Aligned, 2 MisAligned, and 1 Missed outcomes. This indicates that, given only seed transaction(es) from public posts, TxRay supports a triage-and-analysis workflow: it filters non-ACT alerts and produces expert-aligned root-cause reports for 17/20 ACT incidents. The remaining failures mirror the limitations in Section 5.1, pointing to lifecycle reconstruction and ACT-boundary reasoning as bottlenecks.
6.2 Root Cause Analysis Latency Comparison
We compare TxRay against public human responses on incidents that satisfy two conditions: (i) the original alert post does not include a root cause explanation and (ii) the incident is ACT. We mark these 15 cases by bolding the incident names in Table 2. For each case, we collect the quote/reply thread of the original alert post and identify the first response that provides a correct root cause analysis; we record its timestamp as Expert Response. We also record TxRay’s stage completion timestamps (Monitor Finished, Root Cause Finished, and PoC Finished) and compare Expert Response against TxRay’s Root Cause Finished time.
Among the 15 cases, 13 have an expert root cause response. TxRay reaches a root cause earlier than the first expert response in 10 incidents; 8 of these have aligned root causes (DMi, HTC Token, WADJET, PRXVTai, Valinity, AiPay, SwapNet (base), and XPlayer). Experts are earlier in 3 incidents (Metaverse Token, SynapLogic, and Makina Finance), by at most 15 minutes. Across the 15 cases, TxRay’s stage latencies are: postmonitor median 4 minutes (range: 1–9), monitorroot cause median 40 minutes (range: 19–65), and root causePoC median 18.5 minutes (range: 8–53; 14/15 have a completed PoC). This comparison ties end-to-end latency to explicit stages (seed acquisition, root cause analysis, and PoC synthesis) and supports bottleneck analysis and model/tool trade-off studies under a fixed incident stream.
6.3 Ablation Study
To understand which parts of TxRay’s prompt-and-workflow design contribute most to prospective performance, we compare TxRay with two single-agent variants (Table 2, last three column groups). TxRay-Compact is a compact baseline that retains the full downstream tool access (e.g., Foundry, Etherscan, QuickNode, and a decompiler) but uses a simplified single-agent prompt that provides only (i) our ACT definition, (ii) the output requirements and acceptance criteria, and (iii) a complete tool interface. TxRay-Monolith is also single-agent and retains the same full downstream tool access, but it aggregates (iv) distilled role instructions and key constraints from each TxRay sub-agent prompt, (v) the intermediate artifact specifications, and (vi) the handoff requirements between stages, to test the effectiveness of TxRay’s workflow design. Table 2 reports the resulting ablation-study results using the RC abbreviations (AL/MA/MS/NA) and PoC markers (//–). All three settings consistently label the same 4 non-ACT incidents. Starting from the baseline TxRay-Compact, adding explicit workflow structure in TxRay-Monolith yields a clear step up in root cause accuracy: the number of aligned root causes increases from 5 to 11, while misaligned and missed outcomes decrease from 6 and 9 to 5 and 4, respectively. The full TxRay further improves alignment to 17 and reduces errors to 2 misaligned and 1 missed cases. In terms of executable reproductions, correct PoCs increase from 5 (TxRay-Compact) to 10 (TxRay-Monolith) to 17 (TxRay), and all 17 TxRay PoCs are compilable and executable on an on-chain fork. Overall, these results suggest that making the workflow explicit materially improves prospective performance, and the complete TxRay design provides an additional, consistent gain.
7 Generalized Attack Imitation
Although TxRay targets postmortem reconstruction, its oracle-validated executable PoCs can also support attack imitation: given an observed ACT exploit, TxRay outputs a self-contained reproduction that can enable defensive responses (e.g., censoring malicious transactions, recovering victim funds, or, when latency allows, white-hat preemptive value extraction). Prior work includes APE [qin_usenix23], which copies profitable pending transactions and redirects value via dynamic taint analysis and program synthesis, and STING [zhang2023your], which synthesizes counterattack contracts from detected attacks and improves coverage via heuristics for attacker-controlled entities (e.g., contract age, Tornado Cash fund provenance, and source-code verification status). Neither system formalizes a semantic class equivalent to our ACT opportunities (Section 3.2); we evaluate how often TxRay’s pipeline yields executable artifacts that cover historical ACT incidents relative to these baselines. We seek to answer the research question:
RQ8 (Imitation Coverage). How many additional opportunities can be reused for generalized attack imitation by TxRay compared to APE and STING?
| Date | Name | Loss | T. | A. | S. | Date | Name | Loss | T. | A. | S. |
| 21-08 | Popsicle | $20M | 21-08 | Poly Network | $611M | ||||||
| 21-08 | Cream Finance | $18M | 21-09 | DAO Maker | $4M | ||||||
| 21-09 | Nimbus | $5.2K | – | 21-09 | NowSwap | $1.1M | |||||
| 21-09 | Sushi Miso | $3.1M | 21-10 | Indexed Finance | $16M | ||||||
| 21-10 | Cream Finance | $130M | 21-11 | MonoX Finance | $31M | ||||||
| 21-12 | Visor Finance | $8.2M | 22-01 | Anyswap | $1.4M | ||||||
| 22-01 | Qubit Finance | $80M | 22-02 | TecraSpace | $63K | – | |||||
| 22-02 | Build Finance | $470K | 22-03 | Bacon Protocol | $1M | ||||||
| 22-03 | Li.Fi | $570K | 22-03 | Umbrella | $700K | ||||||
| 22-03 | Auctus | $726K | 22-03 | Revest Finance | $11.2M | ||||||
| 22-03 | Ronin Network | $624M | 22-04 | Beanstalk Farms | $182M | ||||||
| 22-04 | Akutar NFT | $34M | 22-04 | Rari Capital | $80M | ||||||
| 22-04 | Saddle Finance | $10M | 22-05 | ApeCoin (APE) | $1.1M | ||||||
| 22-06 | Inverse Finance | $1.2M | 22-06 | SNOOD | $103K | ||||||
| 22-06 | XCarnival | $3.9M | 22-07 | FlippazOne NFT | $1.2K | – | |||||
| 22-07 | Omni NFT | $1.4M | 22-07 | Audius | $1.1M |
Dataset. STING evaluates known Ethereum-mainnet attacks in 2021–2022 [zhang2023your]; APE evaluates 1 Aug 2021–31 Jul 2022 (365 days, inclusive) on Ethereum and BSC and reports its top-100 profitable victim transactions [qin_usenix23]. For a fair comparison, we use the overlap period 1 Aug 2021–31 Jul 2022 and select all Ethereum-mainnet DeFiHackLabs incidents within this window, excluding (i) PoCs or vulnerability demonstrations (Sandbox, Redacted Cartel, CompoundTUSDSweepTokenBypass) and (ii) non-ACT incidents (e.g., private-key compromise; Harmony_multisig). Since this window predates the gpt-5.1 training cutoff (30 Sep 2024) and neither APE nor STING are open-sourced (preventing post-cutoff comparison), we mitigate memorization risk by inspecting all model reasoning traces and verifying that TxRay derives exploits from on-chain evidence (traces, state diffs, and contract interactions). Table 3 summarizes the resulting 32 incidents.
Coverage. Table 3 reports per-incident imitation coverage. TxRay covers all 32 cases and produces an oracle-validated executable PoC for each. STING covers 27/32 incidents (84.4%). APE covers 10/29 evaluated incidents (34.5%) and does not report the remaining 3 below its loss threshold. On this dataset, 5 incidents are only covered by TxRay.
8 DeFi Incident Benchmark
TxRay unifies existing incident datasets into a consistent benchmark with standardized, oracle-validated executable PoCs at low cost, supporting downstream applications such as benchmarking vulnerability detection tools (IDSs), evaluating vulnerability discovery tools (fuzzers, LLM-based exploit generators [gervais2025ai]), and training security models.
DeFi Incident Sources. We aggregate incidents from public incident dashboards and security explorers (DeFiHackLabs, DeFiLlama, De.Fi, Rekt, and BlockSec) and our live incident monitor. We include an entry only if (i) the provider specifies at least one seed transaction hash (and the corresponding chain) and (ii) the incident occurs on one of the 22 TxRay-supported EVM chains; each accepted incident is stored as a standardized artifact bundle with a normalized taxonomy.
Benchmark Statistics. For each provider entry, we feed the reported seed transaction hash(es) to TxRay, which performs data collection, root cause analysis, oracle generation, PoC synthesis, and forked execution validation. We admit an incident to the benchmark snapshot only if the generated PoC both executes successfully and satisfies all generated oracles; otherwise, we exclude it as an incomplete reconstruction. When multiple providers describe the same incident, we run TxRay once and record all source attributions. In the current snapshot, the benchmark covers incidents from DeFiLlama (58), De.Fi (8), Rekt (22), and BlockSec (54) within a 488-day window (1 Oct 2024–31 Jan 2026), as well as the live pipeline (17); to our knowledge, this is the most comprehensive collection of DeFi incidents with executable PoCs within this window. We also include incidents used in our evaluations, including the DeFiHackLabs benchmark (105) and the generalized imitation benchmark (32). In total, the benchmark contains 196 incidents across 9 chains. Each incident is distributed as a Foundry project containing a self-contained, single-file exploit test (test/Exploit.sol) and a command template to reproduce it (e.g., RPC_URL=<...> forge test -vvv).
9 Discussion and Limitations
Dependence on the OpenAI agent stack. Our prototype uses the OpenAI Agents SDK and Codex CLI; we do not evaluate other agent frameworks, tool interfaces, or model providers. Reported latency and cost reflect this stack and the model/pricing configuration at evaluation time.
Scope of ACT incidents. TxRay targets ACT opportunities on EVM-compatible chains only. Attacks involving compromised keys, phishing/social engineering, privileged orderflow, or cross-chain/off-chain components are out of scope.
Complementing human expertise. TxRay achieves strong historical performance and beats expert responses in most prospective cases, but does not yet match the best researchers on all incidents. Traditional analysis remains necessary for complex cases; TxRay is best positioned as a force multiplier for routine postmortems.
Memorization. Throughout this paper, we evaluate TxRay on post-cutoff incidents (after the gpt-5.1 training cutoff, 30 Sep 2024) across the DeFiHackLabs benchmark, live postmortem pipeline, and incident dataset. The sole exception is generalized imitation (Section 7), which uses 1 Aug 2021–31 Jul 2022 because neither APE nor STING have open-sourced their code, preventing a post-cutoff comparison. To mitigate memorization risk, we inspect all reasoning traces to verify that TxRay derives exploits from on-chain evidence via explicit reasoning; readers should interpret imitation coverage results with this caveat.
Generalized Imitation vs. Mempool-time Execution. Section 7 shows that TxRay produces oracle-validated executable PoCs for the generalized imitation dataset, but its end-to-end latency is minutes (Figure 5 (right), Figure 7), far above an Ethereum block interval (12 s) and the same-block reaction budget. Most latency comes from model-facing reasoning (Figure 7), with additional overhead from on-chain data collection and third-party tooling (Figure 8). Online front-running requires either sub-block end-to-end latency or an execution environment that extends the reaction window by delaying inclusion of the adversary transaction. TxRay prioritizes evidence-grounded root-cause reporting and validation; imitation instead requires an executable value-capture strategy, not a human-readable narrative.
Dual use. Automated exploit reproduction is a double-edged sword: it can accelerate both researchers and attackers.
10 Related Work
Incident Investigation / Root Cause Analysis Prior work uses LLMs for incident investigation in cloud operations [chen2023empowering, chen2024automatic] and DFIR/SOC workflows [loumachi2024advancing, loumachi2025advancing, habibzadeh2025large, dai2025automated], motivating verification to mitigate hallucinations [bhandarkar2024digital, cherif2025dfir, ji2024sevenllm]. Blockchain execution provides public ground truth that anyone can verify at scale; TxRay outputs verifiable artifacts by grounding claims in traces/receipts/balance diffs and accepting postmortems only when executable PoCs satisfy semantic oracles.
Smart Contract Security Prior work applies LLMs to smart contract security for vulnerability detection [boi2024smart, yu2025smart, david2023you], economic exploit analysis [gao2025airaclex, liu2025llm], and reviews [he2024large, qian2023empirical], targeting pre-deployment auditing. TxRay focuses on post-incident reconstruction from on-chain evidence using LLM-assisted decompilation for unverified contracts [david2025decompiling].
Executable PoC Generation and Exploit Synthesis Prior work generates executable artifacts from different starting points (reports/findings, templates, or contract-under-test discovery), rather than starting from an on-chain incident transaction. SmartPoC turns textual audit findings into executable Foundry tests validated via differential oracles [chen2025smartpoc]; it targets confirmation of reported bugs rather than reconstructing a specific exploit transaction’s lifecycle. PoCShift migrates existing incident PoCs to new targets by abstracting templates from documented incidents [sunlearning], but requires prior PoCs and targets recurring vulnerable code rather than explaining a concrete incident from chain evidence. Agentic exploit-generation systems iteratively search for profitable exploits on forked states [gervais2025ai], but are designed for vulnerability discovery and output attacker-side exploits without producing evidence-backed postmortems. Given only an on-chain exploit transaction, TxRay (i) reconstructs the lifecycle and root-cause mechanism from public evidence and (ii) synthesizes a deterministic reproduction whose correctness is defined by incident-specific semantic oracles. Unlike audit-centric PoC generators, TxRay treats the incident transaction as the specification: an accepted PoC must reproduce the incident’s on-chain effects and satisfy the oracles. This seed tx grounded root cause oracle-checked PoC pipeline enables standardized, executable benchmarks for real DeFi incidents.
11 Conclusion
TxRay is an agentic postmortem system that reconstructs ACT exploits from seed transaction(s) by collecting public on-chain evidence and producing an evidence-backed root-cause explanation plus a deterministic, self-contained Foundry PoC. It grounds claims in traces/receipts/state diffs and accepts cases only when the PoC satisfies incident-specific semantic oracles, enabling standardized benchmarking and continuous curation of executable incident datasets. We view TxRay as a step toward faster, more reliable DeFi incident response.
Appendix A Walkthrough (PRXVTai Case)
PRXVT is a crypto project that describes itself as privacy-first infrastructure for AI agents, including px402 for privacy-preserving payments. Its ecosystem is associated with a token named PRXVT on Base. We present an end-to-end walkthrough of one incident, aligned with the design in Figure 2.
Iteration summary.
Incident: Base (1 seed tx).
Root-cause: analyzer iterations ; data collection iterations ; fetched items ; challenger: Pass.
PoC: oracle ; reproducer ; validator: Pass.
Seed (09:06 UTC on 1 Jan 2026). CertiK Alert posted:444https://x.com/CertiKAlert/status/2006653156927889666 “Our alert system has detected suspicious transactions involving @PRXVTai…Wallet 0x740 bridged 32.8 ETH from Base to Ethereum and 0x702 still holds ~36.3M PRXVT tokens.” TxRay integrates with the Twitter/X live stream to ingest such posts and extract candidate seed transactions. In this incident, it extracted the Base transaction 0x88610208…5494. TxRay is not an IDS. The choice of seeds and feeds to monitor is external to TxRay.
Bootstrap (09:14 UTC). The orchestrator creates a session directory and an input record (raw.json). It records the seed and fetches on-chain artifacts into the workspace.
Orchestrator input. {"targets":[{"chainid":8453,"txhash":"0x8861…5494"}]}… Workspace writes. raw.json and seed artifacts under artifacts/root_cause/seed/....
After bootstrap, the orchestrator alternates between tool calls and workspace checks, selecting the next tool and prompt based on prior outputs. Because the tools share a workspace, TxRay passes file paths rather than raw data between tools. Below we show excerpted prompts and responses for each turn; Appendix E lists the prompt templates.
Turn 1: Root-cause analyzer (09:19). The orchestrator asks the analyzer to summarize the seed’s effects and emit data_requests. The analyzer identifies a reward-drain call chain and requests evidence to reconstruct the lifecycle.
Prompt (orchestrator Analyzer).
Seed: Base(8453) tx 0x88610208…5494. Task: summarize effects; emit data_requests. …
Response.
Call chain: EOA 0x7407…2f45 -> 0x7029…bce9 -> helper 0xF3FE…410c -> PRXVTStaking::claimReward.
AgentTokenV2 diff: PRXVTStaking -2.297e23, burn +2.297e22, orchestrator +2.067e23. …
Next. Unverified orchestrator/helper; fetch code + tx neighborhood.
Workspace writes. artifacts/root_cause/root_cause_analyzer/iter_0/current_analysis_result.json.
Turn 2: Data collector (09:21). The orchestrator routes data_requests to the collector. Because the contracts are unverified, the collector uses bytecode-level evidence (disassembly + decompilation via Heimdall) and pulls txlists to find related lifecycle transactions.
Prompt (orchestrator Data Collector). Fetch: decompile(0x7029…bce9, 0xF3FE…410c) + txlists (EOA/orchestrator/victims). … Response. Evidence: contract decompiles (no verified source) + txlists with repeated 0xe6d7db7e calls. … Next. Identify minimal lifecycle txs; request traces+difs. Workspace writes. artifacts/root_cause/data_collector/data_collection_summary.json (+ decompiles/txlists under artifacts/root_cause/data_collector/iter_1/...).
Turn 3: Root-cause analyzer (09:25). Given txlists and decompiled code, the analyzer selects lifecycle transactions and requests execution traces and balance diffs. This run expands from one seed into six related transactions spanning setup, reward-drain iterations, and exit.
Prompt (orchestrator Analyzer, refine). Incorporate decompiles/txlists; pick lifecycle txs; request tx_trace+balance_diff. … Response. Requested traces/difs for: 0x7cf175…a5a8 (prepare/stake), 0xe1a6c6…47a8 (early exploit), 0x91d8e0…81f1 (reward-drain). … Workspace updates. Adds the expanded tx set to root_cause.json/all_relevant_txs (final: 6 txs).
Lifecycle mining. Seed expansion identifies lifecycle transactions beyond the seed. Starting from the seed, the analyzer extracts adversary and victim addresses, and the collector pulls account txlists around the seed (explorer API). The analyzer clusters these transactions by counterparty and function selector to identify repeated entrypoints (here, calls to selector 0xe6d7db7e). From this cluster it selects a covering set spanning (i) prepare (stake/receipt setup), (ii) reward-drain, and (iii) exit (profit and principal consolidation). The orchestrator persists this lifecycle set for validation and PoC synthesis.
root_cause.json excerpt:
"all_relevant_txs":[
"0x7cf175…a5a8", "0x886102…5494", "0xe1a6c6…47a8",
"0x91d8e0…81f1", "0x04c182…b7ea", "0x20094a…7d10"] …
Turn 4: Data collector (09:30). The orchestrator fetches the requested traces/diffs (and a mid-sequence transaction) and stores them in the workspace.
Prompt (orchestrator Data Collector). Fetch traces+difs for requested txs (+ mid-sequence 0x04c182…b7ea). … Response. Evidence: prepare stakes ~2.3e24 AgentTokenV2 (mints stPRXVT); exploit transfers the stPRXVT position through helper contracts that call earned()+claimReward(). … Workspace writes. Trace logs and balance diffs under artifacts/root_cause/data_collector/iter_2/....
Turn 5: Root-cause analyzer (09:33). The analyzer identifies the bug and drafts the lifecycle.
Prompt (orchestrator Analyzer, finalize).
Emit root_cause.json grounded in traces/difs. …
Response.
Root cause: transferable stPRXVT does not update reward accounting on transfer (userRewardPerTokenPaid), enabling reward double-counting via helper shuttling.
Lifecycle (excerpt): pre-state B=40230817; prepare 0x7cf175…a5a8 -> reward-drain 0x91d8e0…81f1 -> withdraw 0x20094a…7d10. …
Workspace writes. root_cause.json.
Turn 6: Root-cause challenger (09:44). An independent challenger checks completeness and evidence alignment and either returns structured feedback or accepts; on Pass it also generates the human-readable report from artifacts (sources/bytecode, traces, and diffs).
Prompt (orchestrator Challenger).
Check completeness + evidence alignment; emit Pass/Fail with hints. …
Response.
overall_status: Pass. Checks (excerpt):
-- Checks missing reward-accounting update on stPRXVT transfers in PRXVTStaking.sol.
-- Checks helper-shuttle pattern in traces and decompiled orchestrator/helper.
-- Checks deltas (e.g., tx 0x91d8e0…81f1: pool -2.297e23, orchestrator +2.067e23 AgentTokenV2). …
Workspace writes. artifacts/root_cause/root_cause_challenger/root_cause_challenge_result.json and root_cause_report.md.
Turn 7: Oracle generator (09:47). The oracle generator derives semantic acceptance criteria that check reproduction without replaying attacker-side artifacts.
Prompt (orchestrator Oracle Generator).
Input: root_cause.json. Task: derive semantic hard/soft oracles. Output: oracle_definition.json. …
Response.
Oracles: 3 hard + 2 soft.
Hard: (i) asset identity is AgentTokenV2@0xc2ff…4bc0; (ii) totalStaked unchanged during reward drain;
(iii) a new helper (no stake history) sees earned()>0 and claimReward() yields .
Soft: (i) attacker cluster net AgentTokenV2 increases; (ii) PRXVTStaking reward pool decreases. …
Workspace writes. oracle_definition.json.
These oracles encode the mechanism: if rewards are keyed to current stPRXVT balance (rather than time-weighted ownership), then transferring a matured stake to a new helper can unlock rewards for that helper while leaving principal (totalStaked) unchanged; profit and pool depletion define the ACT success predicate in the reference asset.
Turn 8: PoC reproducer (09:48). The reproducer synthesizes a self-contained Foundry project that forks Base at the incident block and encodes the semantic oracles.
Prompt (orchestrator Reproducer).
Generate a self-contained Foundry PoC on a Base fork at B=40230817; encode oracles as assertions. …
Response (excerpt).
PoC emitted: forge_poc/ with test/Exploit.sol.
Exploit.sol (truncated):
uint256 constant FORK_BLOCK = 40230817;
staking.stake(minimumStake); staking.transfer(orchestrator, stakedBalance);
vm.warp(periodFinish); staking.transfer(helper, orchestratorStake);
vm.prank(helper); staking.claimReward(); …
Workspace writes. forge_poc/ (self-contained project).
Turn 9: PoC validator (09:52). The validator runs forge test and checks oracle satisfaction and the quality rubric.
Prompt (orchestrator PoC Validator).
Run forge tests; check oracle satisfaction + quality rubric. …
Response.
overall_status: Pass.
Correctness: forge test passes on a Base fork (test testExploit_PRXVTStaking_HelperDoubleCounting).
Oracle alignment: all pre_check, hard, and soft constraints are implemented as explicit assertions.
Quality: labeled, self-contained, no attacker identities/artifacts, and no magic numbers (stake derived from minimumStake). …
Workspace writes. poc_report.md and structured validation results.
Timeline. In this run, TxRay performs 17 on-chain fetches and produces a validated root-cause report in 30 min and a validated PoC in 8 min (38 min total from enqueue to PoC validation). This completes within 47 min of CertiK’s initial alert and more than an hour before CertiK’s own root-cause post at 11:13 UTC.555https://x.com/CertiKAlert/status/2006685174587605315 After a manual sanity check, we posted a public root-cause summary on X at 09:58 UTC666URL omitted for double-blind review.
Appendix B Walkthrough (Valinity Case)
Valinity is an Ethereum-based DeFi protocol whose native ERC-20 token is VY. The protocol positions VY as a reserve-backed asset that can be used as collateral to borrow assets from the system. We present an end-to-end walkthrough of a run that requires iterative evidence closure, aligned with the design in Figure 2. For an illustrative case, see Appendix A; this run stresses evidence closure and PoC self-containment, and we mark failure/iteration points in red.
Iteration summary.
Incident: Ethereum (1 seed tx).
Root-cause: analyzer iterations ; data collection iterations ; fetched items ; challenger: Pass.
PoC: oracle ; reproducer ; validator: Pass.
Seed and preflight (02:56 UTC on 4 Jan 2026). The pipeline ingests a single Ethereum seed transaction and confirms that basic on-chain artifacts (metadata/trace/balance diffs) are available in the workspace.
raw.json:
{"targets":[{"chainid":1,"txhash":"0x7f140643…e3395c"}]}
metadata.json (excerpt):
block: 0x17084cf, from: 0xed5a…, to: 0x88F5…, gasPrice: 0x5f5e100.
Seed preflight (excerpt from seed/index.json):
metadata: ok, trace: ok, balance_diff: ok. …
Workspace writes. artifacts/root_cause/seed/index.json and artifacts/root_cause/seed/1/0x7f140643.../.
Turn 1: Root-cause analyzer (iter_0; 03:00). The analyzer extracts strict, trace-grounded facts from the seed and emits targeted data_requests for missing evidence.
Prompt (orchestrator Analyzer).
Seed: Ethereum(1) tx 0x7f140643…e3395c.
Task: summarize effects; emit data_requests. …
Response (excerpt).
Observed: flash-loan-driven swap flow; large ValinityToken mint via a proxy; ETH profit to sender.
Requests: fetch proxy source/ABI; fetch txlists for sender and related addresses. …
Workspace writes. artifacts/root_cause/root_cause_analyzer/iter_0/current_analysis_result.json.
Turn 2: Data collector (iter_1–iter_4; 03:01). Across – iterations (03:01–03:12), the collector incrementally fetches stakeholder code, address histories, and traces. This run already illustrates a common blocker: the attacker-side router is unverified. Compared to analysts with proprietary decompilers and debuggers that pre-process data into high-level views (e.g., DeDaub; BlockSec Phalcon), TxRay uses open-source bytecode tooling (often less accurate) and works directly with low-level RPC/explorer outputs; it therefore grounds feasibility in metadata/txlists and opcode-level evidence (traces and disassembly), and can revise decompiler-derived hypotheses when trace evidence contradicts them.
Prompt (orchestrator Data Collector).
Fetch contract sources/ABIs for Valinity proxy/implementation; pull txlists and traces for candidate roles. …
Response (excerpt from data_collection_summary.json).
Fetched: verified source/ABI for ERC1967 proxy 0x7b4D… and ValinityLoanOfficer 0x8357….
Fetched: address txlists, seed trace (cast run -vvvvv), and opcode-level call traces for a 14-transaction Valinity window around block 0x17084cf (valinity_window_traces_index.json).
Fetched: sigmaB loan parameters (LTV values) into loan_appraisal_summary.json. …
Response (excerpt, router fallback).
Router 0x88F5… is unverified; collected Etherscan metadata + txlist, and later bytecode/disassembly. …
Workspace writes. artifacts/root_cause/data_collector/data_collection_summary.json.
Turn 3: Root-cause analyzer (iter_2, ACT gaps; 03:10). The analyzer quantifies profit and identifies the core mint/loan behavior, but it cannot yet certify ACT feasibility and mechanism without a storage-grounded pre-state reconstruction (sigmaB) and without justifying feasibility when the router is unverified.
What we know (excerpt).
WETH9 loses 22.1188209774 ETH-equivalent;
0xed5a…gains 20.2305043767 ETH and 0x3963…gains 1.8882000143 ETH.
ACT gaps (excerpt).
router 0x88F5... unverified; registrar wiring unknown;
cannot certify unprivileged feasibility without router analysis;
need sigma_B config (caps/LTV/asset registry) at block 0x17084cf. …
Iteration trigger: the seed’s effects are observable, but the ACT claim requires (i) sigmaB configuration (registrar wiring, caps, LTV inputs) and (ii) feasibility checks for an unverified router.
Turn 4: Post-challenge analysis (iter_3, evidence closure; 03:35). To close these gaps, the orchestrator re-invokes the analyzer with explicit evidence-closure requirements (no placeholders; no guessed wiring), and the analyzer emits storage-backed data_requests.
Targeted data_requests (excerpt).
storage_slot: recover registrar and core components at sigma_B.
storage_slot: decode caps and AssetConfig at sigma_B (WBTC/WETH9/PAXG).
state_diff: prestateTracer diffs for the seed tx (key Valinity contracts).
other: analyze router bytecode/disassembly for permissionlessness. …
Turn 5: Post-challenge data collection (iter_5; 03:41). The collector resolves protocol wiring and sigmaB configuration by reading the on-chain registrar and component contracts at the incident pre-state block, and it extracts state diffs across the seed to anchor the mechanism quantitatively.
Prompt (orchestrator Data Collector).
Resolve registrar wiring; read CapOfficer/AssetRegistry state at sigmaB; fetch prestateTracer diffs for the seed tx. …
Response (excerpt from data_collection_summary.json).
Recovered ValinityRegistrar and core components at block 0x17084cf.
Decoded caps/config for WBTC/WETH9/PAXG and captured prestateTracer state diffs. …
Iteration trigger: registrar-based wiring is required to state sigmaB without guessing contract addresses or configs.
Workspace writes. artifacts/root_cause/data_collector/iter_5/contract/1/.../registrar_core_component_mappings_block_0x17084cf.json and artifacts/root_cause/data_collector/iter_5/tx/1/.../state_diff_valinity_prestateTracer_subset.json.
Turn 6: Root-cause finalization and challenge (03:53). With sigmaB grounded via storage reads and feasibility argued via bytecode-level evidence, the analyzer finalizes root_cause.json. The challenger enforces evidence closure (no “unknown” placeholders; no speculative claims); once the report is schema-complete and fully grounded, it accepts and generates a human-readable report.
root_cause.json excerpt (sigmaB).
sigma_B: registrar 0x57DC…maps LoanOfficer, CapOfficer,
AssetRegistry, ReserveTreasury, AcquisitionTreasury, Token.
assets WBTC/WETH9/PAXG: supported and not acquisitionPaused. …
root_cause.json excerpt (b and feasibility).
b: single tx 0x7f1406…e3395c (flash loan + swaps +
VY mint + cap inflation + loans + WETH9ETH profit).
feasibility: router unverified, but disassembly shows no
msg.sender checks; adversary can deploy an equivalent router. …
root_cause.json excerpt (success predicate).
profit (ETH): value_before 18.4981, value_after 40.6168,
delta 22.1187; fees 0.0003166. …
Decision (excerpt).
overall_status: Pass.
‘‘the previously
"unknown
" valuation and verification placeholders have been concretely resolved.’’ …
Workspace writes. artifacts/root_cause/root_cause_challenger/root_cause_challenge_result.json, root_cause.json, root_cause_report.md.
PoC generation: self-containment vs protocol preconditions. The oracle generator specifies semantic acceptance criteria; the reproducer must satisfy them without relying on attacker-side identities or incident-deployed contracts.
Turn 7: Oracle generator (04:02).
Oracle PoC loop (excerpt).
Hard constraints: H1_vy_total_supply_increases, H2_paxg_cap_increases,
H3_paxg_cap_above_collateral_after, H4_vy_collateralized_loan_opened.
Soft constraints: S1_attacker_eth_profit, S2_weth9_reserve_depletion. …
Iteration trigger: a self-contained PoC cannot call the incident router, but the acquisition path depends on on-chain price conditions; the final PoC re-creates these conditions via a USDCPAXG Uniswap V3 swap.
Workspace writes. artifacts/poc/oracle_generator/oracle_definition.json and artifacts/poc/reproducer/reproducer_notes.md.
Turn 8: Reproducer–validator loop (3 iterations; 04:03). This case requires iterative refinement to satisfy both self-containment and feasibility: a “good-enough” human PoC might simply replay incident calldata with hard-coded parameters and attacker-side artifacts, but our validator enforces a higher bar. Accordingly, (i) the validator rejects PoCs that invoke attacker-side artifacts (the incident router); and (ii) even after replacing it with a local router, the exploit requires reproducing the correct on-chain price state for acquireByLTVDisparity() to succeed.
Iteration 1 (04:10, Reject): PoC invoked incident router 0x88F5…; violates self-containment.
Iteration 2 (04:16, Reject): local router introduced, but acquireByLTVDisparity() reverts without price preconditions.
Iteration 3 (04:27, Pass): local router executes USDCPAXG swap (Uniswap V3) to recreate
sigmaB price conditions, then triggers acquisition + loans and satisfies all oracles.
Turn 9: PoC validator (final; 04:27).
Result (excerpt from forge-test.log).
[PASS] testExploit()
Attacker ETH delta: 1888000000000000000
WETH9 ETH delta: -16847219596237316941
VY totalSupply delta: 39645619576378794699219855
PAXG cap before / after: 30075643081474703341424 1275422219460269402561279
Workspace writes. artifacts/poc/poc_validator/poc_validated_result.json and poc_report.md.
Timeline. This run was enqueued by the incident monitor at 02:56 UTC on 4 Jan 2026. TxRay produced a validated root-cause report in ~65 min and a validated PoC in an additional ~27 min (~92 min total from enqueue to PoC validation), including two PoC validation rejections and revision rounds.
Appendix C TxRay Supported Chains
TxRay supports 22 EVM chains: Ethereum (1), Optimism (10), BNB Chain (56), Gnosis (100), Unichain (130), Polygon (137), Monad (143), Sonic (146), zkSync Era (324), HyperEVM (999), Sei (1329), Abstract (2741), Mantle (5000), Base (8453), Arbitrum One (42161), Arbitrum Nova (42170), Celo (42220), Avalanche C Chain (43114), Linea (59144), Berachain (80094), Blast (81457), and Scroll (534352).
Appendix D Incident Monitor Subscribed Accounts
Incident Monitor subscribes to the following Twitter/X accounts: TenArmorAlert, BlockSecTeam, SlowMist_Team, MistTrack_io, BeosinAlert, ChainAegis, MetaSec_xyz, 1nf0s3cpt, PeckShieldAlert, TikkalaResearch, GoPlusZH, exvulsec, blockaid_, getblock_en, WelksCrypto, CertiKAlert, GoPlusSecurity, QuillAudits_AI, 0xfirmanregar, Phalcon_xyz, ACai_sec, audit_911, d23e_AG, De_FiSecurity, hklst4r, MevRefund, nn0b0dyyy, peckshield, AuraAudits, CyversAlerts, hackenclub, OKLink, and SpecterAnalyst.
Appendix E Key Prompt Templates
We show abbreviated templates (with placeholders) for the orchestrator’s tool prompts. Each template is instantiated with session paths and chain-specific values; gray … marks omitted details due to space constraints.
(1) Root cause analyzer (initial).
Inputs: raw.json (seed chainid/txhash), artifacts/root_cause/seed/ (trace, balance diff, receipts/logs), schemas.
Rules: do not browse; do not speculate; every claim must be grounded in provided artifacts.
Task: (i) summarize the seed’s effects (value movement, privileged calls, state changes);
(ii) propose a concrete root-cause hypothesis with the minimal lifecycle (funding, setup, exploit, exit);
(iii) emit explicit data_requests to fetch missing evidence (e.g., tx neighborhood, sources/bytecode,
decompile, traces, balance diffs), each with a destination path under artifacts/root_cause/.
Output (JSON): {summary, hypothesis, candidate_contracts, candidate_roles,
all_relevant_txs (may start with seed), data_requests:[{type, target, reason, out_path}]}. …
(2) On-chain data collector.
Inputs: data_requests (type/target/out_path), chainid_rpc_map.json, .env API keys.
Tooling: QuickNode archive+debug_traceTransaction / logs / storage reads; Etherscan v2
(account.txlist, getsourcecode, getcontractcreation); Foundry (cast/forge);
Heimdall for decompilation when verified sources are unavailable.
Task: satisfy each request deterministically; record failures and fallbacks (e.g., sourcebytecodedecompile).
Outputs: write requested artifacts under artifacts/root_cause/data_collector/iter_k/ and emit
data_collection_summary.json={fetched:[{request, files}], failed:[{request, error}]}. …
(3) Root cause challenger.
Inputs: root_cause.json draft + supporting artifacts (traces, sources/bytecode/decompile, txlists, diffs).
Rules: treat root_cause.json as untrusted; verify against artifacts; reject speculation.
Task: check (i) correctness: each causal step matches on-chain evidence; (ii) completeness: lifecycle txs and roles are covered;
(iii) actionable: the mechanism (violated invariant + code-level breakpoint) is explicit enough to reproduce.
Decision: Accept or Reject with missing_evidence requests and minimal edits.
Output (JSON): {status, feedback, missing_evidence:[{type, target, reason, out_path}]}. …
(4) Poc oracle generator.
Inputs: validated root_cause.json (lifecycle txs, roles, contracts, mechanism, fork block).
Task: translate the root cause into semantic oracles that define successful reproduction without bit-for-bit replay:
hard oracles for critical invariants and permission/ownership changes; soft oracles for profit/loss
with tolerances; also specify prerequisites (fresh roles, required approvals, required funding) and stop conditions.
Output (JSON): {fork_block, hard:[{name, check, target, expected}], soft:[{name, check, tolerance}],
setup:[{action, params}], success_criteria}. …
(5) Poc reproducer.
Inputs: oracle_definition.json, chainid, fork_block, rpc_url.
Rules: the PoC must be self-contained and deterministic: use a fresh EOA(s), do not assume attacker private keys,
and do not depend on external state beyond the fork.
Task: synthesize a Foundry project that reconstructs prerequisites (approvals, funding, role setup) and executes the
attack/MEV flow needed to satisfy the oracles; encode oracles as explicit assert statements and emit useful logs.
Outputs: write forge_poc/ with foundry.toml, test/Exploit.sol, and a short README.md. …
(6) Poc validator.
Inputs: forge_poc/, oracle_definition.json.
Task: run forge test on a fork; confirm each oracle; then score the PoC with the quality rubric
(readability, magic numbers, self-containment, lifecycle documentation).
Decision: Accept if tests pass on a fork without mocked contract behavior and the rubric is computed; else Reject with minimal fixes (e.g., missing setup).
Outputs: write poc_report.md plus poc_validation.json={overall_status, oracle_results, rubric}. …
Appendix F PoCEvaluator Output Schema
"no_real_attacker_side_address": { "description": "Does the PoC avoid using the real attacker-side address, and instead use clean/deterministic Foundry addresses/roles (e.g., makeAddr)?", "evaluation_history": [ { "round": "<int_starting_from_0>", "action": "<Initial|Maintain|Change>", "result": "<true|false>", "reason": "<concise_reason_for_true_or_false>" } ] }
Appendix G DeFiHackLabs Postmortem Outcomes
We show the postmortem outcomes for the 114 DeFiHackLabs incidents in Table 4.
| Date | Incident | ACT | RC | C1 | C2 | C3 | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | Date | Incident | ACT | RC | C1 | C2 | C3 | Q1 | Q2 | Q3 | Q4 | Q5 | Q6 | ||||||||||||||||||
| TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | TR | DH | ||||||||
| 24-10 | FireToken | 25-03 | wKeyDAO | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | LavaLending | 25-03 | DCFToken | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | AIZPTToken | 25-03 | BBXToken | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | HYDT | 25-03 | Alkimiya_io | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | SASHAToken | 25-03 | LeverageSIR | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| 24-10 | P719Token | – | – | – | – | – | – | – | – | – | – | 25-04 | AIRWA | ||||||||||||||||||||||||||||||
| 24-10 | MorphoBlue | 25-04 | Laundromat | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| 24-10 | Erc20transfer | 25-04 | Unverified_6077 | ||||||||||||||||||||||||||||||||||||||||
| 24-10 | VISTA | 25-04 | YBToken | – | – | – | – | – | – | – | – | – | – | ||||||||||||||||||||||||||||||
| 24-10 | CompoundFork | 25-04 | BTNFT | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | RPP | 25-04 | Lifeprotocol | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | ChiSale | 25-04 | ImpermaxV3 | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | CoW | 25-05 | Nalakuvara | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | X319 | 25-05 | MBUToken | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | DeltaPrime | – | – | – | – | – | – | – | – | – | – | 25-05 | Unwarp | ||||||||||||||||||||||||||||||
| 24-11 | vETH | 25-05 | KRCToken_pair | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | MFT | 25-05 | RICE | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | MainnetSettler | 25-05 | UsualMoney | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | Matez | 25-05 | Corkprotocol | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| 24-11 | Ak1111 | 25-06 | unverified_8490 | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | proxy_b7e1 | 25-06 | AAVEBoost | ||||||||||||||||||||||||||||||||||||||||
| 24-11 | NFTG | 25-06 | MetaPool | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | Pledge | 25-06 | BankrollStack | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | LABUBU | 25-06 | BankrollNetwork | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | CloberDEX | 25-06 | Gangsterfinance | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | JHY | 25-06 | GradientMakerPool | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | BTC24H | 25-06 | unverified_b5cb | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | SlurpyCoin | 25-06 | ResupplyFi | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | Moonhacker | 25-06 | Stead | ||||||||||||||||||||||||||||||||||||||||
| 24-12 | Bizness | 25-07 | FPC | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | LAURAToken | 25-07 | unverified_54cd | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | sorraStaking | 25-07 | RANTToken | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | 98Token | 25-07 | gmx | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Mosca | 25-07 | VDS | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | IPC | 25-07 | WETC_Token | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | HORS | 25-07 | Stepp2p | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | LPMine | 25-07 | SWAPPStaking | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | JPulsepot | 25-07 | MulticallWithETH | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | RoulettePotV2 | 25-07 | SuperRare | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Unilend | 25-08 | WXC_Token | – | – | – | – | – | – | – | – | – | – | ||||||||||||||||||||||||||||||
| 25-01 | Mosca2 | 25-08 | Bebop_dex | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | IdolsNFT | 25-08 | YuliAI | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Paribus | 25-08 | coinbase | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | Ast | 25-08 | Grizzifi | ||||||||||||||||||||||||||||||||||||||||
| 25-01 | ODOS | 25-08 | PDZ | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | PeapodsFinance | 25-08 | SizeCredit | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | FourMeme | 25-08 | d3xai | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | unverified_d4f1 | 25-08 | MulticallWithXera | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | StepHeroNFTs | 25-08 | 0x8d2e | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | unverified_35bc | 25-08 | ABCCApp | ||||||||||||||||||||||||||||||||||||||||
| 25-02 | HegicOptions | 25-08 | 0xf340 | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | Pump | 25-08 | EverValueCoin | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | OneInchFusionV1 | – | – | – | – | – | – | – | – | – | – | 25-08 | Hexotic | ||||||||||||||||||||||||||||||
| 25-03 | UNI | 25-09 | Kame | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | SBRToken | 25-10 | MIMSpell3 | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | DUCKVADER | 25-10 | TokenHolder | ||||||||||||||||||||||||||||||||||||||||
| 25-03 | H2O | 25-11 | BalancerV2 | – | – | – | – | – | – | – | – | – | |||||||||||||||||||||||||||||||
| Total | 114 | 109 | 105 | 105 | 114 | 105 | 112 | 105 | 114 | 104 | 109 | 103 | 82 | 53 | 14 | 105 | 3 | 93 | 67 | 105 | 103 | ||||||||||||||||||||||